THE JAHN-TELLER (JT) EFFECT IN THE Ã STATE OF THE NITRATE RADICAL NO_{3}

\underline{KANA TAKEMATSU}, NATHAN EDDINGSAAS, and MITCHIO OKUMURA, Arthur Amos Noyes Laboratory of Chemical Physics, California Institute of Technology, Pasadena, California 91125; JOHN STANTON, Department of Chemistry, University of Texas at Austin, Austin, TX 78712.

The JT effect in the Ã^{2}E′′ of NO_{3} is poorly understood. A preliminary spectrum of the vibronically-allowed Ã←X̃ transition, coupled with ab initio calculations, shows moderate JT activity in the Ã state. Vibronic bands exhibit either static or dynamic JT distortions depending on the vibrational level of the upper Ã state. The picture of the Ã state is however incomplete. For example, in the E′′⊗e′=a_{1}′′⊕a_{2}′′⊕e′′ manifold, while the splitting would provide a direct measure of the JT strength, only the a_{1}′′ levels have been observed. We have gained new insight into the Ã state by examining the hot bands of NO_{3} which access previously unobserved dark levels of the Ã state.